OpenStudy (anonymous):

What is the solution to the proportion 6/10 = 9/x A. 15 B. 5.4 C. 13 D. 30

OpenStudy (bee_see):

15. Mutiply 10 times 9. Gives you 90. Then divide by 6. 90/6=15.

OpenStudy (anonymous):

Thanks could you help me with a few more?

OpenStudy (bee_see):

Sure. Let's see if I can solve them.

OpenStudy (anonymous):

Okay I'l tag you

OpenStudy (anonymous):

B 5.4

OpenStudy (anonymous):

How so? @nathanff

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

15

OpenStudy (anonymous):

x[6/10]=x[9/x] 6x/10=9 10[6x/10]=10[9] 6x=90 6x/6=90/6 x=15